The Chef and I is determining whether it should submit a bid to cater a wedding. In the past, The Chef and I’s main competitor, Beyond Details, has submitted bids 60% of the time.If Beyond Details does not bid on a job, the probability that The Chef and I will get the job is 0.70. If Beyond Details bids on a job, the probability that The Chef and I will get the job is 0.10. (a) If The Chef and I gets the job, what is the probability that Beyond Details did not bid? (Round answer to three decimal places.) (b) What is the probability that The Chef and I will get the job? (Round answer to three decimal places.).
Answer #1

a)

P(not bid | get the job) =P(not bid)*P(get the job|not bid)/P(get the job)

=0.4*0.7/(0.6*0.1+0.4*0.7)=0.824

b)

probability that The Chef and I will get the job =P(not bid)*P(get the job|not bid)+P(bid)*P(get the job|bid)

=0.4*0.7+0.6*0.1 =0.340